# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v2

@@ 6 State Perception
As for scene reconstruction, due to the foundation model’s robust open-set detection capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are then integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. 3D-LLM (Hong et al. (2023)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023)), endowing 3D features with semantic information. CLIP-NeRF (Wang et al. (2022a)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object textures during rendering.
Following Fig. 6, representation learning facilitates the learning of the multimodal encoder, aligning features from various modalities to the text domain, aiding the model’s cognition, akin to how humans learn and think through words. The learning of the encoder promotes open-set perception tasks, such as open-set detection, open-set segmentation, and open-set scene reconstruction. When applied to specific scenarios, models often require fine-tuning. CLIP-Adapter (Gao et al. (2024)) offers a fine-tuning method. However, fine-tuning involves different techniques for different model architectures.
Figure 6: Foundation Models for State Perception. The representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation, and open-set scene reconstruction. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)), and LERF (Kerr et al. (2023)) utilizes the CLIP (Radford et al. (2021)).
@@ 7 Policy
In the realm of manipulation, policies can generate signals for robots’ execution based on input. The types of outputs from policies encompass code, delta pose, and target pose. Policies generate code for robot execution directly, aiding detailed observation by humans for debugging. Generating delta poses directly through policies allows for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose, which often requires subsequent motion planning.
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), and Vision-Language-Action-Delta-Pose (VLADP) based on different output action.
@@ 7.2 VLATP
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image.
@@
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) and BC-Z (Jang et al. (2022)) demonstrate the capability of imitation learning in language-conditioned general manipulation. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023b)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@ 7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)). However, RobotNet can only achieve one-model corresponding to one-task.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024a)) for fine-tuning to construct RFMs is considered efficient. GR-1 (Wu et al. (2023a)) is initially pre-trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates sample data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). Therefore, RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses. Human intervention is added, enabling robots to adjust their trajectories promptly based on human input. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023c)) emerges, effectively enabling a robot manipulation policy with VLMs.
@@
Utilizing pre-trained foundation models without fine-tuning reduces the data collection and training steps. Voxposer (Huang et al. (2023c)) utilizes RGB-D observations of the environment and language instruction as input, then LLMs generate code that interacts with VLMs to produce a sequence of 3D affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation.
@@
Table 1: Robot Foundation Models.This table organizes information about robot foundation models, specifically focusing on robotic manipulation tasks. In the ‘dataset’ section, only manipulation datasets are considered. Four collection methods are defined in ‘Collection Method’: Teleoperation (Human control a robot to perform tasks from a distance, such as the use of VR headsets with controllers), videos (Tasks performed by other robots or humans are recorded for the robot to mimic.), simulation (Tasks are collected in a simulation environment), and skill libraries (Pre-trained models are used to execute in new scene to enhance the dataset with the check module to recognize the task success). ‘Size’ refers to the parameters of the model. ”Input Modality” refers to the modality in which data is inputted into the model. ‘Architecture’ refers to the structure of the model. ‘Hardware’ refers to the devices used during the model training stage. ‘Output modality’ shares the same definition in Section 7. ’Frequency’ refers to the rate at which the model inference and control. ‘Benchmark’ refers to standard tests, focusing solely on manipulation, used to measure the performance.
@@ 18 NVIDIA
60 individual datasets
across 22 embodiments
@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, a plethora of literature investigates the use of foundation models to hierarchically solve challenges like long-horizon problems, effectively exploring and reward function design.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
Following Fig. 7 and Table. 1, pre-trained foundation models can engage with Policy modules in various forms. However, no benchmark indicates which approach is the most effective at present. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
@@ 8 Manipulation Data Generation
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023a)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.

### v3

@@ 6 State
As for scene reconstruction, due to the foundation model’s robust open-set detection capability for objects, DFF (Shen et al. (2023)), CLIP-Fields (Shafiullah et al. (2022)) and LERF (Kerr et al. (2023)) employ CLIP to extract features from multi-view 2D images for NeRF (Mildenhall et al. (2021)) reconstruction. These features are then integrated as part of the output of the NeRF network, enriching the semantic information of the reconstructed 3D scenes. 3D-LLM (Hong et al. (2023b)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023b)), endowing 3D features with semantic information. CLIP-NeRF (Wang et al. (2022a)) integrates semantic features extracted by CLIP into NeRF reconstruction to change object textures during rendering.
Figure 6: Foundation Models for State. The pose estimation method uses the abilities of the VFM encoder, VGM encoder, and VLM language encoder to extract features. These features are integrated into a generative model. The model produces a Normalized Object Coordinate Space (NOCS) map. This map is then post-fitted with the depth map to determine the corresponding pose(Cai et al. (2024)). As for scene reconstruction, using the features extracted by the VLM image encoder as supervision labels for NeRF. This method enables the constructed scene to include semantic information. Using the VLM image encoder to extract features from 3D data multi-view images and lift them into 3D features can incorporate semantic information into the 3D features. The lifting methods include direct reconstruction, gradSLAM, and Neural Field (Hong et al. (2023b)).
Summary
Following Fig. 6, the pose estimation method primarily uses the generative model to generate a Normalized Object Coordinate Space (NOCS) map (Cai et al. (2024)). The scene reconstruction method uses features from the VLM image encoder, such as CLIP (Gao et al. (2024)). These features serve as supervision labels for NeRF. This approach ensures that the reconstructed scene contains semantic information (Kerr et al. (2023)). The VLM image encoder can extract features from multi-view images. These features can then be lifted to 3D space (Hong et al. (2023b)).
@@ 7 Policy
In the realm of manipulation, policies can generate signals for robots’ execution based on input. The types of outputs from policies encompass code, target pose, and delta pose. Policies generate code for robot execution directly, aiding detailed observation by humans for debugging. Generating delta poses directly through policies allows for conversion into trajectory through time sequences, offering a more end-to-end approach compared to target pose, which often requires subsequent motion planning.
The policy model for outputting delta pose resembles more closely the paradigm of human task execution, as it does not require camera and spatial calibration or robot body configuration. Instead, it takes observation images as input and directly outputs the direction and magnitude of the next movement. While this approach is more end-to-end, it still necessitates extensive data training to embed the parameters of robot execution in the policy model’s hidden layers. RT-2 (Brohan et al. (2023)) refers to this approach as Vision-Language-Action (VLA). Following this naming convention, we divide the policy into Vision-Language-Action-Code (VLAC), Vision-Language-Action-Target-Pose (VLATP), Vision-Language-Action-Delta-Pose (VLADP) based on different output action.
@@ 7.2 VLATP
Utilizing foundation models trained on existing large-scale internet datasets enables the direct perception of observation images and outputting target poses. Instruct2Act (Huang et al. (2023b)) utilizes CLIP and SAM to identify manipulated objects within observation images and outputs the 3D position of these manipulated objects from 2D image. DALL-E-Bot (Kapelyukh et al. (2023)) employs DALL-E to generate target images for tasks and generates target poses for manipulation by combining the target image with the observation image. ReKep (Huang et al. (2024)) uses VFM and VLM to extract relational keypoint constraints from language instructions and RGBD observations. It then applies an optimization solver to generate a series of end-effector poses.
@@
As for imitation learning methods, CLIPort (Shridhar et al. (2021)) demonstrates the capability of imitation learning in language-conditioned general manipulation. However, CLIPort (Shridhar et al. (2021)) addresses 4-DoF end-effector pose prediction by treating it as a pixel classification problem. Keypoint-based approaches are extended to handle 6-DoF end-effector pose prediction. Due to keypoint-based methods primarily focus on 3D scene-to-action tasks, these methods become computationally expensive as resolution requirements increase (Ke et al. (2024)). To address high spatial resolution, PerAct (Shridhar et al. (2023)) uses the latent set self-attention of Perceiver (Jaegle et al. (2021)), which has linear complexity with voxels. Act3D (Gervet et al. (2023)) represents scenes as a continuous 3D feature field, transforming 2D model features into 3D feature clouds using sensed depth. It learns 3D features through recurrent coarse-to-fine point sampling.
Some research has extended the work on PerAct (Shridhar et al. (2023)) and Act3D (Gervet et al. (2023)). ChainedDiffuser (Xian et al. (2023)) builds upon Act3D (Gervet et al. (2023)) by replacing the motion planner with a diffusion model. This approach addresses the challenges of continuous interaction tasks. The 3D Diffuser Actor (Ke et al. (2024)), similar to Act3D (Gervet et al. (2023)), employs tokenized 3D scene representations. However, unlike Act3D (Gervet et al. (2023)) and 3D Diffusion Policy (Ze et al. (2024b)) with 1D point cloud embeddings, 3D Diffuser Actor (Ke et al. (2024)) leverages CLIP to extract features from 2D images and aggregates them into a 3D scene representation. GNFactor (Ze et al. (2023)) improves upon PerAct (Shridhar et al. (2023)) by enhancing 3D semantic features. It achieves this by distilling pre-trained semantic features from 2D foundation models into Neural Radiance Fields (NeRFs). DNAct (Yan et al. (2024)) builds on GNFactor (Ze et al. (2023)) and transforms the perceiver model into a diffusion head. VoxAct-B (Liu et al. (2024b)) uses VLM to divide the task into subtasks for the left arm and the right arm and applies PerAct (Shridhar et al. (2023)) to generate separate target poses for each arm.
RFMs based on the above work are also under development. LEO (Huang et al. (2023a)) expands upon language foundation models by incorporating modalities like images and 3D point clouds. It fine-tunes manipulation datasets using the LoRA method. This showcases the ability to transfer the original foundation model to more modalities and manipulation tasks. VIMA (Jiang et al. (2023)) and MIDAS (Li et al. (2023b)) observe that many robot manipulation tasks can be represented as multimodal prompts intertwining language and image/video frames. They construct multimodal prompts manipulation datasets and utilize pre-trained language foundation models for fine-tuning to control robot outputs. Xu et al. (2024) considers the task goal, the object’s physical properties, and the end-effector’s design and creates a ManiFoundation model to generate the target pose. However, the target pose output by the ManiFoundation model is not 6D pose. Instead, it provides the positions of multiple contact points and the force to be applied at each contact point.
@@ 7.3 VLADP
RoboNet (Dasari et al. (2019)) establishes an initial pool of 15 million video frames from 7 different robot platforms, aiming to learn a generalizable model for vision-based robotic manipulation through an end-to-end approach. This learning method closely mimics human learning, eliminating the need for calibration and 2D-to-3D mapping to understand spatial 3D coordinates. Due to imagination, RoboNet generates predicted 2D images of the next movement and estimates which pixel the robot should move to in the next step, ultimately converting it into the action the robot should take. This concept of imagination is applied in both 3D-VLA (Zhen et al. (2024)) and SuSIE (Black et al. (2023)). However, RobotNet can only achieve one-model corresponding to one-task. One-model corresponding to one-task methods have also achieved significant progress. ACT (Zhao et al. (2023a)) uses action chunks to reduce compounding errors. Diffusion policy (Chi et al. (2023)) applies the idea of diffusion to visuo-motor control, tackling challenges such as action multimodality and sequential correlation to handle high-dimensional action sequences.
@@
Utilizing pre-trained VLMs (Zhang et al. (2024a)) for fine-tuning to construct RFMs is considered efficient. RT-2 (Brohan et al. (2023)) collects manipulation trajectory data and fine-tunes manipulation datasets using VLM models like PaLI-X (Chen et al. (2023b)) and PaLM-E (Driess et al. (2023)) after treating delta pose as tokens. However, this approach necessitates a lot of data for the hidden layers to learn parameters related to the robot body, objects, and environment. Open X-Embodiment (Padalkar et al. (2023b)) assembles a dataset from 22 different robots, demonstrating 527 skills (160266 tasks). However, the current Open X-Embodiment dataset faces the heterogeneity dataset challenge. Octo (Team et al. (2024)), RDT-1B (Liu et al. (2024d)), HPT (Wang et al. (2024c)) and PI (Black et al. (2024)) propose multi-module networks to address this issue. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction. However, democratizing such an expensive framework for all robotics practitioners proves challenging as it relies on private models and necessitates extensive co-fine-tuning on vision-language data to fully exhibit effectiveness. Consequently, there is an urgent need within robot communities for a low-cost alternative solution, hence RoboFlamingo (Li et al. (2023c)) and OpenVLA (Kim et al. (2024)) emerge, effectively enabling a robot manipulation policy with VLMs.
Internet videos contain information on the physics and dynamics of the world, some studies have explored training VLMs using both video datasets and manipulation data. GR-1 (Wu et al. (2023a)) is initially trained on a large-scale video dataset for video prediction, and then seamlessly fine-tuned with manipulation data. GR-2 (Cheang et al. (2024)) uses VQGAN to convert each image into discrete tokens and is trained with a larger text-video dataset than GR-1. LAPA (Ye et al. (2024)) begins by extracting the latent delta action between video frames. It then labels the video dataset with this information. These labeled datasets are used to train a VLM network. Finally, a small-scale robot manipulation dataset is applied for fine-tuning, enabling the mapping of latent delta actions to robot actions. At the same time, BC-Z (Jang et al. (2022)) and Vid2Robot (Jain et al. (2024)) introduce a video-conditioned policy that uses video as task instructions. MUTEX (Shah et al. (2023)) extends instruction to various modalities and develops speech-conditioned, speech-goal-conditioned, image-goal-conditioned, and text-goal-conditioned.
Previous studies, such as GR-1 (Wu et al. (2023a)) and GR-2 (Cheang et al. (2024)), train the policy head using MSE regression. In contrast, OpenVLA (Kim et al. (2024)) and RT-2 (Brohan et al. (2023)) apply classification loss for their policy head. Building on the success of diffusion policy (Chi et al. (2023)), TinyVLA (Wen et al. (2024b)) adopts a diffusion head as its policy head, achieving better performance than OpenVLA.
@@
Utilizing pre-trained foundation models without fine-tuning reduces the data collection and training steps. Voxposer (Huang et al. (2023c)) utilizes LLMs generate code that interacts with VLMs to produce value maps affordance maps and constraint maps, collectively referred to as value maps, grounded in the robot’s observation space. These composed value maps serve as objective functions for motion planners to synthesize trajectories for robot manipulation.
@@
Table 1: Robotic-specific Foundation Models.This table organizes information about robotic-specific foundation models, specifically focusing on robotic manipulation tasks. In the ‘dataset’ section, only manipulation datasets are considered. Four collection methods are defined in ‘Collection Method’: Teleoperation (Human control a robot to perform tasks from a distance, such as the use of VR headsets with controllers), videos (Tasks performed by other robots or humans are recorded for the robot to mimic.), simulation (Tasks are collected in a simulation environment), and skill libraries (Pre-trained models are used to execute in new scene to enhance the dataset with the check module to recognize the task success). ‘Size’ refers to the parameters of the model. ”Input Modality” refers to the modality in which data is inputted into the model. ‘Architecture’ refers to the structure of the model. ‘Hardware’ refers to the devices used during the model training stage. ‘Output modality’ shares the same definition in Section 7. ’Frequency’ refers to the rate at which the model inference and control. ‘Benchmark’ refers to standard tests, focusing solely on manipulation, used to measure the performance.
@@ 18 NVIDIA
Open X-Embodiment
(Padalkar et al. (2023a))
@@ 10 NVIDIA
OpenVLA
(Kim et al. (2024))
Open X-Embodiment
(Padalkar et al. (2023a))
DROID
(Khazatsky et al. (2024))
Human Demonstrations1
Teleoperation
7B
Image
Language
Llama 2 7B
64 NVIDIA
A100 GPUs
Delta Pose 6HZ
BridgeData V2
(Walke et al. (2023))
RT-1, RT-2
(Brohan et al. (2022, 2023))
3D Diffuser Actor
(Ke et al. (2024))
RLBench
(James et al. (2020))
CALVIN
(Mees et al. (2022))
Simulation
Teleoperation
RGBD
Language
Diffusion Model
Target Pose 1.67HZ
RLBench
(James et al. (2020))
CALVIN
(Mees et al. (2022))
DNAct
(Yan et al. (2024))
RLBench
(James et al. (2020))
Simulation
3D Point Cloud
Language
Diffusion Model
2 NVIDIA
RTX 3090 GPUs
Target Pose -
RLBench
(James et al. (2020))
@@ 7.4 Foundation Models assisting for Reinforcement Learning
Reinforcement learning has garnered widespread attention from researchers due to its ability to explore the environment by not requiring extensive annotated data. However, it also faces numerous challenges, such as dealing with long-horizon sequences, effectively exploring, reusing experience data, and designing reward functions (Kober et al. (2013)). Foundation models have demonstrated the emergence of common sense reasoning, the ability to propose and sequence sub-goals, and visual understanding. Due to the strong capability of foundation models, many studies aim to leverage the unprecedented capability of foundation models to address the challenges faced by reinforcement learning. RobotGPT (Jin et al. (2024)) aims to distill the knowledge of the brain ChatGPT into the mind of a small brain trained with reinforcement learning. At the same time, many studies explore the use of foundation models to solve challenges like long-horizon problems and effectively exploring and designing reward functions.
Norman (Di Palo et al. (2023)) employs LLMs to decompose tasks into subgoals and utilizes CLIP to identify the completion of each subgoal, serving as a signal generator for sparse rewards. Eureka (Ma et al. (2023b)) utilizes LLM to craft a reward function for five-fingered hand pen spinning. Subsequently, it engages in a cyclic process encompassing reward sampling, GPU-accelerated reward evaluation, and reward reflection to progressively refine its reward outputs. In contrast to Eureka’s self-iteration and sparse reward function design, TEXT2REWARD (Xie et al. (2023a)) incorporates human feedback into the iterative updating of the reward function, yielding a dense reward function. FAC (Ye et al. (2023a)) proposes using knowledge from foundation models as policy prior knowledge to improve sampling action efficiency, as value prior knowledge to measure the values of states and as success-reward prior knowledge to provide final feedback on task success.
Summary
Following Fig. 7 and Tab. 1, pre-trained foundation models can engage with Policy modules in various forms. However, no benchmark indicates which approach is the most effective at present. Regarding end-to-end RFMs, there are still many architectural considerations to make models more interpretable and reliably stable.
@@ 8 Manipulation Data Generation
Generating lots of data in simulation is a cheaper solution. However, it still requires human effort to create both scene generation and task execution code for specific tasks (Wang et al. (2023b)). Moreover, the notorious sim-to-real gap issue remains a challenge in transferring policies trained in simulation to real-world applications. But there are many methods to address the sim-to-real challenge. Matas et al. (2018) trains the policy fully in simulation through domain randomization and then successfully deployed in the real world, even though it has never encountered real deformable objects. Therefore, simulation plays an important role in manipulation and this section will analyze existing simulators, scene generation, demonstration generation and sim-to-real gap challenge. Regardless of whether it’s in a real or simulated environment, improving the efficiency of the existing dataset is essential. The mainstream approach is dataset augmentation.